Identify the accounts receivable section in the 2018 Amazon annual report. After reviewing this information, explain how accounts receivable can impact liquidity and solvency ratios. Additionally, you may notice some companies use total receivables instead of net receivables. Discuss how this distinction might impact the calculation and liquidity analysis. Provide an example and describe how bad debt affects liquidity and solvency ratios.
Answer #1

16677 million is the account receivable in 2018 Amazon annual report.

Liquidity ratios are financial analysis tools commonly used to gauge a company's ability to repay short-term creditors out of its cash fund. Liquidity ratios measure a company’s liquid assets against its short-term liabilities.

The solvency ratio is a key metric used to measure an enterprise’s ability to meet its debt obligations .The solvency ratio is one of many metrics used to determine whether a company can stay solvent. Solvency ratios include debt-to-equity, total-debt-to-total-assets, and interest coverage ratios.

Accounts receivable can impact liquidity ratios but there is no impact on solvency ratios. As solvency ratios are for long term effects , and accounts receivable is Current Asset used for assessing Short term impact. More the accounts receivable , better will be the current assets and higher will be the liquidity ratio. Companies with low liquidity ratios have a higher risk of encountering difficulty meeting obligations.

Net receivable = Total receivable - Bad debt

Total receivables can be used in companies where there are very less chances of bad debt. Companies use net receivables to measure the effectiveness of their collections process and to make projections of anticipated cash flows.

Net receivables is a more conservative principle and works on Prudence basis. If estimated bad debt is not subtracted from total receivables, inflated liquidity ratios can be reported, which will be inaccurate for any decision making. Solvency ratio will also get affected as bad debt affects the profit figure and consequent tax liability.

Solvency Ratio= Net After − Tax Income + Non-Cash Expenses​ / Short-Term Liabilities + Long-Term Liabilities

For eq - Total Receivable is 100

bad debt is 10

current liability is 50

Total liabilty is 150 (short term and long term)

Income is 20 and tax is nil

LIQUIDITY RATIO with TOTAL RECEIVABLE = 100/50 = 2

LIQUIDITY RATIO with NET RECEIVABLE = 100-10 / 50 = 1.8

SOLVENCY RATIO with TOTAL RECEIVABLE= 20 / 150 = 0.13

SOLVENCY RATIO with NET RECEIVABLE = 20-10 / 150 = 0.067